ROTATIONAL SPECTROSCOPY OF ZnCCH (X^{2}Σ^{+})AT MICROWAVE AND MILLIMETER WAVE-LENGTHES

J. MIN, D. T. HALFEN, M. SUN, B. T. HARRIS, L. M. ZIURYS, University of Arizona, Deptment of Chemistry and Biochemitry and Steward Observatory, Tucson, AZ-85721.

The pure rotational spectrum of ZnCCH (X^{2}Σ^{+}) has been measured using Fourier transform microwave (FTMW) and direct absorption millimeter/submillimeter methods in the frequency range of 7-260 GHz. This is the first study of ZnCCH by any spectroscopic technique. In the FTMW system, the molecule was synthesized using discharge assisted laser ablation spectroscopy (DALAS) from a mixture of 0.05% acetylene in argon and the ablation of a zinc rod. In the millimeter-wave spectrometer, the radical was created from the reaction of zinc vapor, produced in a Broida-type oven, with HCCH in a DC discharge. Spectra of the main isoplogue, ^{64}ZnCCH, as well as ^{66}ZnCCH, ^{68}ZnCCH, ZnCCD and Zn^{13}C^{13}CH have been recorded. The data have been analyzed with a ^{2}Σ Hamiltonian and rotational, spin-rotation and H, D and ^{13}C hyperfine parameters have been determined. The structure is being calculated based on the rotational constants and will be presented. Interpretation of the hyperfine constants will also be discussed.